OpenStudy (anonymous):

does anyone know how to rearrange equations? like: p = w/t (solve for w & t)

OpenStudy (anonymous):

\[p=\frac{w}{t} \]To solve for w, multiply each side of the equation above by t. The t's on RHS divide out, t/t = 1, leaving the following:\[p t= w \]Now divide each side of the last result above by p. The p's on the LHS divide out, p/p =1 leaving:\[t=\frac{w}{p} \]The following is a list of the results in Traditional form:\[\left\{p=\frac{w}{t},w=p* t,t=\frac{w}{p}\right\} \]